OpenStudy (anonymous):

Subtract: 18.4 – (- 18.9) is it -37.3

Vocaloid (vocaloid):

nope when we have two negative signs together, it becomes a positive sign 18.4 - (-18.9) = 18.4 + 18.9

OpenStudy (anonymous):

37.3

Vocaloid (vocaloid):

yes

OpenStudy (anonymous):

Multiply: -18 x 2.4

Vocaloid (vocaloid):

negative*positive = negative

OpenStudy (anonymous):

sorry was not supposed send that again so do i make the -18 to 18

OpenStudy (anonymous):

-43.2

Vocaloid (vocaloid):

if it's supposed to be positive 18, then your answer should also be positive

Vocaloid (vocaloid):

18*2.4 = 43.2 -18*2.4 = -43.2 does that make it a bit more clear?

OpenStudy (anonymous):

so a neg and neg becomes a positive right

Vocaloid (vocaloid):

yeah

OpenStudy (anonymous):

Divide: -9 divided by (- 4) (show as an improper fraction

OpenStudy (anonymous):

so 9/4

Vocaloid (vocaloid):

yes

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

a positive and a negitive is a negitive

OpenStudy (anonymous):

Divide: 3.8 divided by (-4) (show as a decimal)

Vocaloid (vocaloid):

just do 3.8/4 on a calculator, then stick a negative sign in front

OpenStudy (anonymous):

Find the reciprocal of the following number: -21/23

Vocaloid (vocaloid):

oh sorry for the late reply just switch the numerator and denominator keep the negative sign.